a pure monopolist faces ademand function of p=20-Q,and the cost function given by TC=2+4Q2.what are the profit maximizing level of output and the corresponding profit respectively?


1
Expert's answer
2021-04-29T10:30:54-0400

A net monopoly maximizes its profit under the condition that:

MC=MR

MC=TC'=8Q

"TR=PQ=(20-Q)Q=20Q-Q^2"

MR=TR'=20-2Q

8Q=20-2Q

Q=2

P=20-Q=20-2=18

"corresponding profit=TR-TC=P\times Q-(2+4Q^2)=18\times2-(2+4\times2^2)=4"